Grade 11General Physics

a ball is dropped from the top of a building, the ball takes 0.5 seconds to fall passs the 3 m lenghth of a window some distance top of the building. with what speed does the ball pass the top of a window?

vatsala
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
Give h a chance to be the separation of the window from where the ball is dropped. 
 
What's more, given in the inquiry is the length of window that is 3 m 
 
Give t1 a chance to be the time taken by the ball to cross window. 
 
Give v1 a chance to be the speed of ball 
 
Lets utilize the newtons first eqn of movement 
 
v2 = v1 + gt1 
 
Furthermore, t1 = 0.5 
 
v2 – v1 =5 - (1.) 
 
and after that utilization the third eqn 
 
v2^2 = v1^2 + 2gl 
 
As g = 10m/s^2 , l = 3m 
 
V2^2 –v1^2 = 60 
 
(v2+v1)(v2-v1) =60 
 
(v2+v1)* 5 = 60 
 
(v2+v1) = 12 - (2.) 
 
By settling eq (1.) and (2.) 
 
V2 = 8.5 m/sec 
 
V1 = 3.5 m/sec 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)